OpenStudy (anonymous):

finding the slope of an equation of a line y=-3x+6

OpenStudy (anonymous):

This is already in the form of y = mx + b where 'm' is the slope. So what's the slope?

OpenStudy (anonymous):

What is the general equation of the straight line

OpenStudy (anonymous):

-3 ?

OpenStudy (anonymous):

Yes

OpenStudy (marinedude42):

thats right, i agree

OpenStudy (anonymous):

help with another?

OpenStudy (anonymous):

Sure

OpenStudy (marinedude42):

why not

OpenStudy (anonymous):

9x-11y=9

OpenStudy (anonymous):

This is not in the form of y = mx + b, so you must change it, by first subtracting 9x to both sides.

OpenStudy (anonymous):

So you have -11y = -9x + 9 Now you have to divide all terms by -11, to get 'y' by itself. y = 9/11x + 9/11

OpenStudy (anonymous):

So what's the slope? @Kaylaaa

OpenStudy (anonymous):

9/11 ?

OpenStudy (anonymous):

YES